Suppose there are two independent economic factors, M1 and M2. The risk-free rate is 7%. Portfolios A and B are both well diversified.

Portfolio Beta on M1 Beta on M2 Expected Return

A 1.8 2.1 40%

B 2.0 -0.5 10%

What is the risk premium for M1?
Answer #1
Expected return = risk free rate+M1 beta*M1 risk premium+M2 beta*M2 risk premium
40=7+1.8*M1 risk premium + 2.1*M2 risk premium
33=1.8*M1 risk premium + 2.1*M2 risk premium……(1)
10=7+2*M1 risk premium + -0.5*M2 risk premium
3=2*M1 risk premium + -0.5*M2 risk premium……(2)
Solving (1) & (2) simultaneously we get
M1 risk premium=4.47